# DEAR: Dataset for Evaluating the Aesthetics of Rendering

### v3

@@ 5 Rendering preferences prediction
We adopt the following notation: the DMOS value for an image pair $i$ and $j$ within scene $s$ is denoted as $p^{s}_{ij}$, with the corresponding predicted score given by $\hat{p}^{s}_{ij}$. The aggregated pairwise score obtained via the Bradley–Terry procedure [bradley1952rank] is denoted as $q^{s}_{i}$, with its predicted counterpart $\hat{q}^{s}_{i}$.
@@ 5.1 Aesthetic preference prediction
In the first block in Table 2 there are results for evaluating total variation (TV) and pretrained models: BRISQUE[Mittal2012], PIQE [pique], ARNIQA [arniqa] CLIP IQA[Wang2023], LARIQA [lariqa] and LIQE [Zhang2023]. It can be seen that simple TV performs comparable with foundation-based models like LIQE. This means that current IQA models are not applicable in an EAR task without finetuning on this data.
@@ 5.2 Personalized EAR
In the previous section, we analyzed the aggregated preferences of all voters. In this experiment, we focus on a single voter: the most active one. This individual voter provided 7,678 votes, covering 595 scenes in which he/she voted on every available style at least once. Given the relatively small size of this subset, we restricted our evaluation to simple models based on handcrafted image features and CLIP embeddings. The results are reported in Table 3. It can be seen that individual prediction performs worse than similar methods in Table 2 except for top-1 accuracy. This may be the result of a well-known issue, where individuals are confident about the best and the worst examples and are indifferent to anything in between [parducci1965category, tversky1974judgment].
@@ 7 Conclusions
8 Supplementary for DEAR: Dataset for Evaluating the Aesthetics of Rendering
8.1 BLIP
To study the dependence of the importance of sliders on the content of the scene, we marked up scenes from MIT-Adobe Fivek using bootstrapping language-image pre-training (BLIP) from Salesforce [^2] https://huggingface.co/Salesforce/blip-vqa-base. The markup prompts were as follows:
  • People: Is it some people in a casual setting?
  • Portrait: Is it close-up photo of a person’s face?
  • Landscape: Is it natural scenery: mountains, fields, water, desert, canyon?
  • Buildings architecture: Is it exterior view of buildings, architectural structures?
  • Interior: Is it indoor room, interior items?
  • Warm interior: Is it indoor scene with warm lighting?
  • Cloudy Outdoor: Is it overcast outdoor scene with diffuse lighting and soft shadows?
  • Midday sunny outdoor: Is it bright outdoor scene with harsh shadows and high contrast?
  • Sunrise sunset: Is it sky with vibrant colors near the horizon during dawn, dusk?
  • Flowers: Is it close-up of blooming flowers or floral arrangements?
  • Animals: Is it photo of an animal: pet, mammal, bird, fish, insect?
  • Food: Is it appetizing close-up of prepared meal?
  • Night: Is it low-light scene taken after dark with artificial lighting?
We got “yes” or “no” answers for each rendering of a scene and then selected the final answer for a scene by the majority of votes. For the accuracy estimation, we manually labeled 100 scenes. The computed metrics are reported in Table 5.
Table 5: Accuracy and other metrics for the BLIP markup.
Accuracy F1 Score Precision Recall TP,% FP,% FN,% TN,%
People 0.92 0.87 0.84 0.90 26 5 3 66
Portrait 0.97 0.90 0.93 0.88 14 1 2 83
Landscape 0.95 0.78 1.00 0.64 9 0 5 86
Buildings architecture 0.96 0.95 0.92 0.97 36 3 1 60
Interior 0.99 0.98 1.00 0.96 25 0 1 74
Warm interior 0.96 0.90 0.82 1.00 18 4 0 78
Cloudy Outdoor 0.91 0.91 0.85 0.98 44 8 1 47
Midday sunny outdoor 0.88 0.57 0.89 0.42 8 1 11 80
Sunrise sunset 0.98 0.93 0.87 1.00 13 2 0 85
Flowers 0.99 0.95 1.00 0.90 9 0 1 90
Animals 0.93 0.70 0.73 0.67 8 3 4 85
Food 0.99 0.94 1.00 0.89 8 0 1 91
Night 0.93 0.84 0.72 1.00 18 7 0 75
8.2 Expert markup analysis
In Figure 9 we report the estimated number of settings used in MIT-Adobe FiveK. Despite having a large variety of sliders to change, approximately 90, the renderers only changed a slight amount of them.
Figure 9: Number of changes for most used settings in MIT-Adobe Fivek.
8.3 Examples of model prediction
In Figure 10 we illustrate some examples of scenes where significantly different levels of aesthetic of rendering are correctly estimated with our model MEAR, whereas other state-of-the-art models (i.e. LIQE and ARNIQA) judge them to be more or less aesthetically equivalent.
Figure 10: Examples of images. Our model shows visibly more attention to rendering preferences than existing methods.

### v4

@@ 5 Rendering preferences prediction
We adopt the following notation: the DMOS value for an image pair $i$ and $j$ within scene $s$ is denoted as $p^{s}_{ij}$, with the corresponding predicted score given by $\hat{p}^{s}_{ij}$. The aggregated pairwise score obtained via the Bradley–Terry procedure [8] is denoted as $q^{s}_{i}$, with its predicted counterpart $\hat{q}^{s}_{i}$.
@@ 5.1 Aesthetic preference prediction
In the first block in Table 2 there are results for evaluating total variation (TV) and pretrained models: BRISQUE[5], PIQE [34], ARNIQA [1] CLIP IQA[37], LARIQA [6] and LIQE [45]. It can be seen that simple TV performs comparable with foundation-based models like LIQE. This means that current IQA models are not applicable in an EAR task without finetuning on this data.
@@ 5.2 Personalized EAR
In the previous section, we analyzed the aggregated preferences of all voters. In this experiment, we focus on a single voter: the most active one. This individual voter provided 7,678 votes, covering 595 scenes in which he/she voted on every available style at least once. Given the relatively small size of this subset, we restricted our evaluation to simple models based on handcrafted image features and CLIP embeddings. The results are reported in Table 3. It can be seen that individual prediction performs worse than similar methods in Table 2 except for top-1 accuracy. This may be the result of a well-known issue, where individuals are confident about the best and the worst examples and are indifferent to anything in between [25, 33].